OpenStudy (swissgirl):

Let R be the relation on \( \mathbb{Z} \) given by \( \{ (x,y):x^2+y=2 \} \) Prove that R is function with domain \( \mathbb{Z} \)

OpenStudy (helder_edwin):

Z is the integers \(\mathbb{Z}\), right?

OpenStudy (swissgirl):

yaa i just dont the symbol of it in latex

OpenStudy (helder_edwin):

just type \mathbb{Z}

OpenStudy (helder_edwin):

well u have \[ \large y=2-x^2 \] so for a given x you can find (compute) y

OpenStudy (helder_edwin):

now suppose \((x,y_1)\) and \((x,y_2)\) are both in R then \[ \large y_1=2-x^2 \] and \[ \large y_2=2-x^2 \] but this imples \(y_1=y_2\) so R is a function

OpenStudy (swissgirl):

what abt proving the component that its domain is in Z?

OpenStudy (helder_edwin):

\[ \large y=2-x^2 \] has no restrictions of any kind (division by zero, roots or logarightms) so its domain is \(\mathbb{Z}\)

OpenStudy (swissgirl):

alrighty. Thanks :DDD

OpenStudy (helder_edwin):

u r welcome